A 6​-ft vertical post casts a 12​-in shadow at the same time a nearby cell phone tower casts a 116​-ft shadow. How tall is the cell phone​ tower?
Answer by jim_thompson5910(35256)


12 inches = 1 ft

Let
A = height of post = 6 ft
B = length of post's shadow = 1 ft
C = height of cell phone tower = unknown
D = length of cell phone tower's shadow = 116 ft

The right triangles formed by the sun's rays are similar triangles, so we can set up the proportion below, plug in the given values, and solve for C.
A/B = C/D
6/1 = C/116
6*116 = 1*C .... cross multiply
696 = C
C = 696

The cell phone tower is 696 feet tall.